# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

### v2

@@ 3.3.1 MolEdit & MolOpt Evaluation
For the MolEdit and MolOpt tasks, as shown in Table 1, the following metrics are adopted:
Success Rate: We have developed automated testing processes with RDKit, which compare the generated molecule with the original molecule and check if the number of the specified functional group is correct after modification (for MolEdit task) and if the specified chemical property is optimized as required (for MolOpt task). Details of the automated testing process are shown in Appendix E. The success rate reflects the percentage of molecules that have successfully passed the rigorous testing processes. However, as it is impossible to track every change between the generated molecules and the original molecule (See Appendix E for detailed explanation), successfully passing the test processes does not necessarily mean that the generated molecule is modified from the original molecule.
Similarity: In this case, we need to evaluate the similarity between the original molecule and the generated molecule. When removing a specified functional group, the modification should follow the minimal change principle. Therefore, a high similarity indicates that the generated molecule can be modified from the original molecule via as few as possible steps, instead of transforming the molecule into an entirely new one or inadvertently removing additional components. For implementation, we calculate the Tanimoto Similarity between the generated molecule and the original molecule by transforming them into the Morgan Fingerprints (Butina, 1999), which can be described by the following equation:
@@
where $m^{g}$ denotes the generated molecule, $m^{o}$ is the original molecule, and $\delta(m^{g},m^{o})$ is the Tanimoto Similarity between the generated molecule and the original molecule, while $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints. $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
Then, the similarity score for a subtask $t$ in MolEdit and MolOpt, denoted as $\delta_{t}$, is determined by taking the average of the similarity scores from all individual test samples.
Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax. We use RDKit to parse the SMILES strings. Only the molecules that successfully pass the parser are considered valid molecules. A higher validity means that the model is more familiar with the grammar of molecule textual representations.
@@ 3.3.2 MolCustom Evaluation
For the MolCustom task, the metrics adopted are also shown in Table 1. As LLMs are required to generate the molecule from scratch, we employ a different testing process for the evaluation of success rate and introduce the novelty score instead of similarity:
Success Rate: For different subtasks in MolCustom, we have designed different testing processes by the chemical toolbox. We calculate the number of the specified atoms, functional groups, or bonds for the three subtasks to check if the generation meets the requirements. Detailed testing process of MolCustom could also be found in Appendix E. A high success rate indicates that the LLM is effectively generating molecules according to the request.
Novelty: However, since there may be numerous valid solutions, fulfilling the basic requirements alone does not guarantee the quality of generation. To enable the exploration of new molecule structures, we employ novelty as an indication of the quality, which compares the similarity between the newly generated molecules and existing ones in the database. The novelty score is particularly important for discovering potential drug or material candidates, where innovation can lead to significant advancements. For implementation, we choose the Zinc250K as the database and calculate the average Tanimoto Similarity between the generated molecules and the existing molecules. The novelty $n$ for the generated molecule $m^{g}$ can be calculated as follow:
$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc250K}\delta(m^{g},m^{k})}{\left|Zinc250K\right|},$ (2)
Then, the novelty score forthe subtask $t$ in MolCustom, denoted as $n_{t}$, is determined by taking the average of the novelty scores from all individual test samples.
Validity: Similarly, validity is also an important metric for the MolCustom task.
@@ 3.3.3 Average Weighted Success Rate
To comprehensively evaluate the average performance of LLMs on TOMG-Bench, we introduce an average weighted success rate to rank the performance of LLMs. Considering that the success rate alone can not ensure the accurate generation for MolEdit and MolOpt tasks and the novelty score for MolCustom is also a critical metric for the quality of molecule generation, we adopt similarity and novelty scores as weights of the success rate:
@@
where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of performance that considers both the success rate and quality of the generation.
@@ 3.4 OpenMolIns: Instruction Tuning Dataset
In this section, we introduce OpenMolIns, a specialized dataset derived from the PubChem database to help LLMs get familiar with text-based open molecule generation via instruction tuning. This instruction tuning dataset is meticulously designed to ensure that the molecules it contains do not overlap with those in the Zinc-250K dataset to promote the generation of more novel molecular structures and avoid any potential data leakage that could compromise the integrity of the model performance.
We collect the instruction tuning dataset by constructing the samples of the nine subtasks in equal amounts. The RDKit toolbox is applied to construct the instruction tuning dataset. Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
For the MolEdit and MolOpt domains, we randomly select a functional group from the original molecule for addition or removal. Then, similarly, we wrap the original molecule and the edited molecule with a random pre-defined prompt to construct the training sample for MolEdit. In the case of MolOpt, we also evaluate the direction of the desired property changes by calculating the properties to determine whether the operation improves or decreases the property value. Then, the original molecule, the optimized molecule, and the optimization direction are altogether wrapped by a pre-defined prompt.
In the MolCustom domain, constructing the instruction tuning samples is rather straightforward. We calculate the molecular statistics and encapsulate them within prompts. For example, for AtomNum, we count all the atoms we would consider in the molecule, including their types and numbers. Then, we wrap these statistics with a random pre-defined prompt to construct the training sample. This method allows the generated molecules to better fit the real distribution of molecular space.
To investigate the impact of data scales on the performance of LLMs, we have established five distinct data levels: light, small, medium, large, and xlarge, as illustrated in Section 3.5. Each level represents a different quantity of data, shown in Table 2, allowing researchers to analyze how the amount of training data influences the model’s ability to generate and edit molecules effectively.
Dataset TOMG-Bench OpenMolIns
@@
Table 2: Statisics of TOMG-Bench and OpenMolIns.
@@ 3.5 Statistics
In this section, we outline the data sizes of TOMG-Bench along with the OpenMolIns dataset, which are shown in Table 2. For TOMG-Bench, we have three main tasks with nine subtasks in total, where each subtask contains 5,000 test samples.
OpenMolIns has five distinct data scales: light, small, medium, large, and xlarge, ranging from 4,500 to 1,200,000 examples, which helps us investigate the data scaling law of applying LLMs to the text-based open molecule generation task. The nine subtasks within the TOMG-Bench are uniformly distributed in OpenMolIns.
4 Experiments
4.1 Models
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful language understanding and generation capabilities. These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
Open-source General LLMs. This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications. Specifically, we benchmark Llama-3-70B-Instruct, Llama-3-8B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), and chatglm-9B (GLM et al., 2024). These LLMs are top-tier open source general LLMs within our computation budget.
Open-source ChEBI-20 Fine-tuned LLMs. LLMs fine-tuned on the ChEBI-20 dataset can grasp some extent of text-based molecule generation capability. In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large (Edwards et al., 2022), and BioT5-base (Pei et al., 2023). These models have been the state-of-art models in the molecule-caption translation task.
OpenMolIns Fine-tuned LLMs We further adopt LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama-3.1-8B-Instruct on OpenMolIns dataset for instruction tuning. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M. We select Galatica-125M because it has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023). Meanwhile, a small size model can also help us study the data scaling law within reasonable budget. Besides, Llama3.2-1B-Instruct and Llama-3.1-8B-Instruct are selected due to their advanced capabilities and similar architectures.
Figure 3: The performance of models benchmarked in TOMG-Bench. In TOMG-Bench, LLMs are divided into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models of unknown parameters are denoted as horizontal lines.
4.2 Findings
Figure 3 presents the performances and model sizes of different models benchmarked on TOMG-Bench, as well as the instruction-tuning performance of Galactica-125M on the five distinct data levels of OpenMolIns. More precise details for each subtask are further illustrated in Appendix F. Based on the above results, we observe the following key findings:
Text-based open molecule generation is both promising and challenging for LLMs. As illustrated in Figure 3, we have calculated the average weighted success rate across all the nine subtasks. Among the LLMs benchmarked, Claude-3.5 stands out as the top performer, achieving an average weighted success rate of 35.92%. Gemini-1.5-pro follows closely with an average weighted success rate of 34.80%. Meanwhile, fine-tuned on OpenMolIns-large, Llama-3.1-8B-Instruct achieves an average weighted success rate of 27.22%, outperforming all the open-source general LLMs.
However, these results still underscore the considerable scope for improvement, even among the most advanced proprietary LLMs. What’s more, despite LLMs like Claude-3.5 and GPT-4o exhibit relatively strong performance in the MolEdit and MolOpt tasks, the more challenging MolCustom task still remains a challenge. In MolCustom tasks, no LLM has managed to achieve a success rate exceeding 25% for a single subtask. Meanwhile, as shown in Figure 5, among the three subtasks of MolCustom, BondNum is the most difficult task, as it requires LLMs to understand the connections or edges in the molecule graph merely from the textual representation. These observations indicate that the generation of molecules from scratch demands a deep understanding of the molecular structures, an area where current models are still striving to make significant strides.
Most powerful open-source general LLMs can already outperform GPT-3.5-turbo. In TOMG-Bench, Llama-3-70B-Instruct has achieved an impressive average weighted success rate of 23.93%, notably outperforming GPT-3.5-turbo, which scored 18.58%. Compared to the proprietary models, open-source general LLMs seem to have been lack of chemical corpora for the pre-training process. However, recent development of open-source general LLMs have rapidly bridged the gap. The evolution of the Llama series, in particular, has been remarkable, finally surpassing GPT-3.5-turbo and showing the increasing diversity of pre-training corpora.
More powerful LLMs usually inherit a better performance in TOMG-Bench. Across all the LLMs we benchmarked, a clear trend emerged: the more powerful the LLM is, the higher performance it can achieve on the TOMG-Bench. For instance, the GPT series has consistently demonstrated improved performance with each new iteration from GPT-3.5-turbo to GPT-4o.
Similarly, within the Llama-3 series, we could also observe that larger and more powerful models tend to achieve superior results on the TOMG-Bench. These findings underscore a strong correlation between an LLM’s capabilities and its performance in text-based open molecule generation.
However, we encountered an unexpected anomaly with certain open-source LLMs. Notably, Qwen2-7B-Instruct, despite its impressive ability to solve mathematical problems and its size of 7 billion parameters, underperformed models with as few as 1 billion parameters. This result is particularly striking and suggests that the TOMG-Bench offers a unique and comprehensive evaluation that current races for LLMs may not have adequately addressed. This discovery also highlights the significance of the TOMG-Bench as a new benchmark for LLMs. It provides a broader and more diverse assessment that exposes potential blind spots in the development of LLMs.
ChEBI-20 is insufficient for LLMs to master molecular structures and editing operations. The ChEBI-20 dataset is designed to bridge the gap between molecules and textual descriptions. Despite this intention, LLMs trained on ChEBI-20 have demonstrated limited effectiveness in TOMG-Bench. For instance, BioT5-base, which is claimed as the state-of-the-art model for text-based molecule generation on the ChEBI-20 dataset, only achieves an average weighted success rate of 4.21% on the TOMG-Bench. In the MolEdit and MolOpt tasks, although these models obtain a high success rate, the generations are actually not modified from the original molecules, leading to poor similarity scores. This also shows that the average weighted success rate provides a more accurate reflection of the actual performance. Similarly, in the MolCustom task, which closely mirrors the text-based molecule generation task, the performance remains unsatisfactory, with no model achieving a success rate above 5% in a single subtask. This performance shortfall highlights a critical limitation of the ChEBI-20 dataset, as it lacks the data quantity and diversity necessary to effectively align molecules with textual descriptions.
The data scaling law varies among different tasks. Across five distinct data scales of OpenMolIns and the experiments on Galactica-125M, we observe a general data scaling law: as the size of the corpus increases, the performance of LLMs also improves. As shown in Figure 3 and Table 12, the outcomes were remarkable: Galactica-125M achieved a weighted average score of 25.73% on OpenMolIns-xlarge, surpassing even the 70B Llama-3-70B-Instruct and GPT-3.5-turbo, with only 125 million parameters.
However, an interesting observation shows that the data scaling law varies among different tasks. For instance, in MolCustom tasks, we witnessed a substantial average improvement of 286% when we scaling the data size from large to xlarge, which suggests that, for specific tasks, increasing data size has a beneficial effect on model performance.
Conversely, in the MolCustom tasks, the performance gains from data scaling were notably modest. More concerning, in the MolEdit tasks, we even observed a decline in performance with increased data size, which indicates that the model’s capacity is not solely data-driven but is also constrained by the number of parameters. It suggests that for certain subtasks within TOMG-Bench, simply increasing the volume of data may not lead to proportional improvements in performance, and other factors such as model complexity or optimization strategies may need to be considered.
5 Conclusion
In this study, we introduce TOMG-Bench, the first benchmark designed to assess the capabilities of LLMs in the realm of open-domain molecule generation. Benchmarking 25 LLMs, TOMG-Bench highlights the limitations of existing targeted molecule generation tasks and demonstrates the potential of general LLMs in this domain. Additionally, we propose an instruction tuning dataset, OpenMolIns, to fully explore the potential of LLMs on the TOMG-Bench, enabling Llama-3.1-8B-Instruct to outperform GPT-3.5-turbo. Our contributions not only lie in the development of a novel benchmark for LLMs in molecular discovery but also provide a diverse indication of the LLM evaluation.
Appendix A Comparison with Existing Text-based Molecule Benchmarks or Datasets
@@
Mol-Instructions (description guided molecule design) (Fang et al., 2024) Targeted Generation 297,319 1000
TOMG-Bench & OpenMolIns Open Generation 1,200,000 45,000
@@
Table 3: Comparison with existing text-based molecule generation benchmarks or datasets.
Table 3 shows the comparison between TOMG-Bench & OpenMolIns and the existing text-based molecule generation benchmarks or datasets. TOMG-Bench introduces the brand-new open-ended molecule generation task for LLMs, which boosts the diversity of text-based molecule generation task. Compared to the existing benchmarks or datasets, OpenMolIns has the largest training set, which could better aligns the molecular space with the natural language space. Meanwhile, TOMG-Bench owns three major tasks and nine subtasks with 45,000 test samples in total, whose test data size is also the largest one among all the existing benchmarks and datasets, enabling more accurate evaluation and comprehensive assessment of LLMs’ performance in text-based molecule generation.

### v3

@@ 2.3.1 MolEdit & MolOpt Evaluation
For the MolEdit and MolOpt tasks, as shown in Table 1, we employ a combination of metrics to assess both the correctness of the modification and the rationality of the design.
Success Rate: This metric evaluates a model’s ability to fulfill the specific molecule generation requirements. We design automated evaluation processes to verify if the generated molecule meets the specified criteria (e.g., correct functional group modification for MolEdit, or desired property optimization for MolOpt). Detailed Implementations of the automated testing process are shown in Appendix E. This metric examines whether the LLM can follow instructions precisely, which is a fundamental requirement for an LLM to serve as a chemist assistant.
Similarity: In open-domain tasks like MolEdit and MolOpt, a high Success Rate alone is insufficient. We must also ensure that the generated molecule is a reasonable modification of the original, rather than a completely new, unrelated structure that happens to satisfy the criteria. To address this, we measure the Tanimoto Similarity between the generated and original molecules using Morgan Fingerprints (Butina, 1999). A high similarity score indicates that the model has performed a rational, localized edit and has not simply generated a different molecule from scratch. The similarity $\delta(m^{g},m^{o})$ between the generated molecule $m^{g}$ and the original molecule $m^{o}$ is calculated as:
@@
where $fp_{m^{g}}$ and $fp_{m^{o}}$ represents their corresponding Morgan Fingerprints. $\left|fp_{m^{g}}\cap fp_{m^{o}}\right|$ is the size of the intersection between their Morgan Fingerprints, while $\left|fp_{m^{g}}\cup fp_{m^{o}}\right|$ is the union.
Validity: This metric ensures that the generated molecules are chemically valid and follow the rules of molecular syntax. Only the molecules that successfully pass the SMILES parser are considered valid molecules. A higher validity means that the model is more familiar with the molecule syntax.
@@ 2.3.2 MolCustom Evaluation
For the MolCustom task, where models are required to generate molecules from scratch based on structural descriptions, we use a different set of metrics to evaluate their creative design capabilities.
Success Rate: This metric measures how well the generated molecules adhere to the high-level structural constraints provided in the prompt (e.g., number of atoms, bonds, or specific functional groups). As the task requires de novo generation, the success rate directly reflects a model’s ability to translate abstract requirements into a valid, concrete molecular structure. Detailed testing process of MolCustom could also be found in Appendix E.
Novelty: For open-domain generation, fulfilling the requirements is just the first step. The true value lies in a model’s ability to generate novel and innovative molecules. The novelty score quantifies this by comparing the generated molecules against existing structures in a large database, such as Zinc-250K (Zinc for short). A low similarity to known molecules suggests a high degree of novelty, which is a critical indicator of a model’s potential for discovering truly new molecule structures. The novelty $n$ for the generated molecule $m^{g}$ can be calculated as:
$n(m^{g})=1-\frac{\sum_{m^{k}\in Zinc}\delta(m^{g},m^{k})}{\left|Zinc\right|},$ (2)
Validity: As with the other two tasks, validity ensures that all generated molecules are chemically sound and can be interpreted correctly.
@@ 2.3.3 Average Weighted Success Rate
To provide a single, comprehensive ranking of LLM performance on S2-Bench, we introduce a weighted success rate. This metric combines the core success rate with a quality metric relevant to each task: Similarity for MolEdit/MolOpt and Novelty for MolCustom. This approach ensures that a high score reflects not only the ability to follow instructions but also the rationality and creativity of the generation. The weighted success rate for a subtask $t$ is defined as:
@@
where the average weighted success rate $\overline{W\!S\!R}$ provides a balanced measure of a model’s performance across all nine subtasks, offering a robust and nuanced assessment of its capabilities in open-domain natural language-driven molecule generation.
@@ 2.4 OpenMolIns: Instruction Tuning Dataset
To effectively train and evaluate LLMs on the open-ended challenges posed by S2-Bench, we introduce OpenMolIns, a specialized instruction-tuning dataset derived from the PubChem database. The core philosophy behind OpenMolIns is to provide models with the kind of flexible, one-to-many training examples they need to learn genuine chemical reasoning, rather than simply memorizing specific input-output pairs. To ensure the integrity of our evaluations, we meticulously designed OpenMolIns to have zero overlap with the Zinc-250K, preventing any data leakage that could compromise the validity of our benchmark results.
We built the instruction-tuning dataset programmatically, creating equal numbers of samples (i.e., 5,000) for all nine subtasks. The RDKit toolbox was essential to this process, allowing us to automatically generate data based on objective molecular properties and rules. This contrasts sharply with previous methods that rely on scarce and costly human annotations. Details of the construction process and the pre-defined prompt templates are shown in Appendix D.
For MolEdit and MolOpt, our goal is to improve the capabilities of LLMs to perform rational, goal-oriented modifications. For MolEdit, we start with a molecule, programmatically identify a modifiable functional group, and then perform a change (add, delete, or substitute) to create a new, valid molecule. This pair of (original and modified) molecules is then used to construct an instruction. This process teaches the model not only a single correct answer, but rather the skill of modifying a molecule in a chemically sound way. Similarly, for MolOpt, we calculate properties before and after a modification to identify a specific optimization direction (e.g., increasing LogP or decreasing MR). The training sample then explicitly links a desired outcome to a structural change, teaching the LLM to reason about the relationship between structure and property.
In the MolCustom domain, we aim to train the model to generate molecules from scratch based on structural details. Instead of relying on existing molecule-caption pairs, we programmatically analyze molecules from PubChem, extract key structural statistics (e.g., atom counts, bond numbers, and functional group types), and use these as the basis for a training prompt. The molecule itself becomes the target output. This approach trains the model to synthesize a new molecule that fits a set of high-level, multi-faceted criteria, a skill fundamentally different from recalling a specific, pre-existing structure. This method allows the generated molecules to better fit the real distribution of molecular space and prepares the model for de novo design.
To further investigate the impact of data scales on an LLM’s ability to learn these tasks, we created five distinct data levels: light, small, medium, large, and xlarge, shown in Table 2. This tiered structure allows researchers to systematically analyze the data scaling law for open-domain natural language-driven molecule generation, providing valuable insights into how data quantity influences a model’s capacity for chemical reasoning.
Dataset S2-Bench OpenMolIns
@@
Table 2: Statisics of S2-Bench and OpenMolIns.
@@ 2.5 Statistics
In this section, we provide a detailed overview of S2-Bench and OpenMolIns, as summarized in Tables 2 and 3. The design and scale of our datasets are not arbitrary; they are meticulously crafted to enable a rigorous and nuanced evaluation of LLMs in open-domain molecule generation, which previous, smaller datasets cannot provide.
S2-Bench is composed of nine subtasks, with each containing 5,000 test samples, for a total of 45,000 carefully curated test cases. This extensive size is critical for providing a statistically robust and comprehensive assessment of model performance, mitigating the risk of misleading results that can arise from smaller test sets. More importantly, these samples are designed to test the model’s ability to handle one-to-many relationships, a challenge that is central to true chemical reasoning.
Meanwhile, to support the development of models capable of succeeding on this benchmark, OpenMolIns is provided in five distinct data scales, ranging from 4,500 to 1,200,000 examples. This tiered structure serves a dual purpose: it not only offers ample data to train powerful models but also allows researchers to systematically investigate the data scaling law for this novel task. By observing how model performance improves with increasing data size, we can gain valuable insights into the data requirements for aligning molecular space with natural language.
@@
Mol-Instructions (Fang et al., 2024) Targeted Generation 297,319 1000
L + M - 24 (Edwards et al., 2024) Targeted Generation 160,492 21,839
S2-Bench & OpenMolIns Open Generation 1,200,000 45,000
@@
Table 3: Comparison with existing text-based molecule generation benchmarks and datasets. Notably, we denote the statistics of the description guided molecule design part in Mol-Instructions.
Besides, as shown in Table 3, S2-Bench and OpenMolIns stand out from existing text-based molecule generation datasets. S2-Bench is the first to introduce a truly open-domain natural language-driven molecule generation task, moving the field beyond the limitations of one-to-one mapping. In terms of data scale, our datasets are also significantly larger. OpenMolIns has the largest training set to date, with a volume that is unprecedented in this domain. This scale is crucial for enabling LLMs to learn complex chemical rules and generalize beyond the memorization of specific examples. Similarly, S2-Bench’s 45,000 test samples make it the largest evaluation benchmark, providing a more reliable and comprehensive measure of a model’s true capabilities. This combination of innovative task design and large-scale data represents a new paradigm for evaluating and advancing LLMs in the field of text-guided molecular discovery.
3 Experiments
3.1 Models
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful reasoning and generalization capabilities. These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
Open-source General LLMs. This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications. Specifically, we benchmark Llama3-70B-Instruct, Llama3-8B-Instruct, Llama3.1-8B-Instruct, Llama3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), chatglm-9B (GLM et al., 2024), and Gemma3-12B (Team et al., 2025). These LLMs are top-tier open-source general LLMs within our computation budget.
Open-source ChEBI-20 Fine-tuned LLMs. LLMs fine-tuned on the ChEBI-20 dataset can grasp some extent of text-based molecule generation capability. In this case, our experiments also cover LLMs like MolT5-small, MolT5-base, MolT5-large (Edwards et al., 2022), and BioT5-base (Pei et al., 2023). These models have been the state-of-art models in the molecule-caption translation task. Here, to ensure a fair evaluation of the intrinsic capabilities of LLMs, we exclude models that utilize multi-modal architectures or retrieval-augmented generation.
OpenMolIns Fine-tuned LLMs. We further fine-tune LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama3.1-8B-Instruct on OpenMolIns dataset for comparison. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M because the model has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023). Meanwhile, a small size model can also help us study the data scaling law within reasonable budget. Furthermore, Llama3.2-1B-Instruct and Llama3.1-8B-Instruct are also selected due to their advanced capabilities and similar architectures.
Figure 2: The performance of LLMs benchmarked in S2-Bench. LLMs fall into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models of unknown parameters are denoted as horizontal lines.
3.2 Findings
Based on our comprehensive benchmarking, we have made the following key observations regarding the capabilities of LLMs in open-domain natural language-driven molecule generation.
F1: Current LLMs show promise but lack the structural understanding necessary for precise molecule generation. As shown in Figure 2, top-tier proprietary models such as Claude-3.5 and Gemini-1.5-Pro attain 35.92% and 34.80% in average weighted success rate, respectively. These results suggest that general LLMs hold promise for open-ended molecule generation due to their powerful reasoning capability, yet substantial challenges remain in translating textual requirements into chemically valid and structurally precise molecules. Notably, within the MolCustom task, no LLM manages to surpass a 25% weighted success rate on any individual subtask. The particularly poor performance on MolCustom tasks highlights that while current LLMs can generate chemically plausible molecules in a broad sense, they struggle with satisfying fine-grained structural constraints such as precise atom, bond, or functional group counts, because the precise control of molecular structures requires reasoning over discrete, globally coupled constraints. LLMs, even trained on SMILES, primarily learn local token distributions from large-scale pre-training data, which rarely emphasizes exact numeric constraints.
The precise control, in fact, is crucial for molecular discovery, as many properties of interest, such as solubility, toxicity, and active site composition, depend on specific structural configurations. Our finding suggests that open-domain natural language–driven molecule generation task can serve as an effective diagnostic tool, revealing systematic weaknesses in current LLMs and providing rich failure cases that can inform future instruction tuning or reinforcement learning strategies to improve molecule structural constraint adherence.
F2: While pre-training provides a foundation for LLM performance, instruction tuning is indispensable for guiding and optimizing their capabilities in molecular discovery. Open-source general LLMs can often lack chemical corpora during their pre-training stage. As a result, models like Qwen2-7B-Instruct and yi-1.5-9B obtained unsatisfactory results. The observed shortcoming can be attributed to the lack of chemical data in pre-training, despite their notable proficiency in mathematics and general language understanding. In contrast, Llama3.1-8B-Instruct delivers superior performance at a relatively modest model scale, demonstrating a more balanced and comprehensive capability across diverse domains.
Meanwhile, our experiments with OpenMolIns highlight the effectiveness of instruction tuning. Fine-tuned on the largest OpenMolIns scale (i.e., xlarge), Llama3.1-8B-Instruct surpassed all other LLMs with an average weighted success rates of 39.33%. More remarkably, the much smaller model, Galactica-125M, within just 125 million parameters, achieved a weighted average success rate of 25.73%, outperforming models two orders of magnitude larger, such as Llama3-70B-Instruct. This remarkable outcome indicates that task-specific instruction tuning on a large, high-quality corpus can be highly effective, suggesting a practical avenue for LLMs to efficiently narrow the performance gap in molecular discovery.
F3: One-to-one mappings between language instructions and molecules make LLMs inherently rely on pattern recognition and recall. Our results also substantiate the observation, demonstrating that existing one-to-one mappings are insufficient for evaluating true chemical comprehension of LLMs. Notably, ChEBI-20 Fine-tuned LLMs, which are designed to align molecules with texts, perform worse than general LLMs. For instance, BioT5-base (Pei et al., 2023), a state-of-the-art model on the ChEBI-20 dataset, achieves an average weighted success rate of only 4.21% on S2-Bench.
A closer examination of BioT5-base’s performance provides insight into this limitation. Although the model achieves relatively high success rates on MolEdit and MolOpt tasks, its generated molecules often exhibit low similarity to the original structures. This indicates that the model is not performing rational edits; rather, it generates different but valid molecules that happen to satisfy the criteria. In other words, BioT5-base relies more on recalling patterns from its training corpus than on learning to systematically modify molecular structures. Consequently, while the ChEBI-20 dataset is valuable for molecule–caption translation, its limited diversity, scale, and one-to-one mapping nature are insufficient for training LLMs to perform nuanced, open-ended molecular generation.
Figure 3: Task-specific performance scaling with increasing data in S2-Bench. (a) Galactica-125M. (b) Llama-3.1-8B-Instruct.
F4: The data scaling law does not always hold: task-specific bottlenecks limit gains in molecule optimization and editing. As shown in Figure 3, our experiments with Galactica-125M across five data scales of OpenMolIns provide new insights into the data scaling law for molecule generation. Overall, we observe that larger datasets generally improve performance, but the effect is highly task-dependent.
In MolCustom, scaling the dataset from large to xlarge yields a remarkable 286% average performance gain, indicating that tasks involving complex, de novo synthesis strongly benefit from larger and more diverse training data. By contrast, Galactica-125M obtains only modest gains in MolOpt, and even negative gains in MolEdit, suggesting that additional data alone cannot overcome existing bottlenecks in these relatively simpler tasks.
Interestingly, when examining Llama-3.1-8B-Instruct, we find that performance continues to increase when scaling from the large to xlarge data level, except for LogP and MR subtasks. This contrast generally highlights that small models like Galactica-125M are capacity-limited, whereas larger models can more effectively leverage additional data.
To summarize, these findings suggest that while complex, open-ended tasks are data-driven, simpler property optimization or editing tasks may be constrained more by model capacity than by dataset size. Thus, merely enlarging training data may not yield proportional benefits. Instead, future progress will require a balanced consideration of both model scale and data scale.
4 Conclusion
In this study, we introduced S2-Bench and OpenMolIns, the first benchmark and instruction tuning dataset for evaluating the capabilities of LLMs in open-domain natural language-driven molecule generation. By moving beyond traditional one-to-one text-to-molecule mappings, our benchmark accommodates the open-ended nature of molecular discovery. S2-Bench focuses on the realistic molecular reasoning and design, rather than simple pattern matching and memorization. Our comprehensive benchmarking of 28 LLMs not only highlights the challenges faced by current models and the limitations of existing targeted molecule generation datasets, but also demonstrates the substantial potential of LLMs for natural language–driven molecule discovery.
Appendix A Related Work
Molecule discovery is a cornerstone of scientific progress, underpinning advances in both drug development and material design (Du et al., 2022). The integration of artificial intelligence into this process has driven a paradigm shift in the pharmaceutical landscape, substantially improving the efficiency and accuracy of identifying and developing novel therapeutic candidates.
Recent breakthroughs in natural language processing (NLP) have highlighted the potential of Large Language Models (LLMs) to analyze complex biological and chemical data more effectively than traditional computational methods (Zhou et al., 2023). Within this context, Text-based Molecule Generation has emerged as a representative task for AI-driven molecule discovery (Edwards et al., 2021). This task focuses on generating target molecules directly from natural language descriptions, which requires the construction of paired datasets linking molecular structures with their textual representations.
Early work in this direction employed transformer-based models such as MolT5 (Edwards et al., 2022), which leveraged large-scale self-supervised learning to produce high-quality SMILES strings from textual prompts. Building on this foundation, models like KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023) integrated molecular graphs and biochemical texts to jointly enhance molecular understanding and generation. More recently, 3D-MoLM (Li et al., 2024d) advanced the field by incorporating spatial configurations, thereby improving the geometric validity and accuracy of generated molecular structures.
Parallel to these architecture-driven improvements, LLMs such as MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) have demonstrated strong in-context learning capabilities, enabling adaptive molecule generation by retrieving and utilizing relevant examples from the provided context. Besides, MolReFlect (Li et al., 2024c), emphasizes fine-grained alignment between molecular structures and natural language descriptions through a teacher–student paradigm, which effectively captures subtle structural–semantic relationships. Moreover, Mol-R1 (Li et al., 2025) takes a step further by introducing DeepSeek-R1 like Long-CoT reasoning in molecule generation, enabling robust and interpretable natural language-driven molecule generation.